# Aligning with Human Judgement: The Role of Pairwise Preference in Large Language Model Evaluators

### v2

@@ 5.1 Experimental Setup
Datasets. We evaluate PairS using three meta-evaluation benchmarks for two types of NLG tasks: abstractive summarization and creative story generation. Following previous research, we adopt two evaluation strategies, namely sample-level and dataset-level Spearman correlations, as detailed in Appendix  §D.1. For summarization task, we select News Room (Grusky et al., 2018) and SummEval (Fabbri et al., 2021), consistent with prior studies (Zhong et al., 2022; Fu et al., 2023). For the story generation task, we include HANNA (Chhun et al., 2022). [^3] We make the first application of pairwise preference to evaluate over thousands samples. We use an anchor size of 100 on PairS-beam-scaled in HANNA. We refer to Appendix §C for more details.
@@
Baselines. We consider two types of reference-free baselines: optimized and zero-shot evaluators. Optimized evaluators, UniEval (Zhong et al., 2022) and BARTscore (Yuan et al., 2021), mitigate the evaluation standard gaps using annotated data, making them specialized for a specific dataset or domain but not transferable. We treat them as strong but unfair baselines and use their reported scores. For zero-shot evaluators, we re-implement the following baselines: BERTScore (Zhang et al., 2020), G-Eval (Liu et al., 2023a), and GPTScore (Fu et al., 2023). We report direct scoring as the main baseline for comparison.
@@
Table 1: LLM evaluation results. We report Spearman correlations on all datasets in terms of aspects: coherence (CH), fluency (FLU), relevancy (RE), informativeness (IN), consistency (CON), surprise (SU), and complexity (CX). For PairS-beam, the experiments were conducted with a beam size of 1000 and uncertainty threshold at 0.6.
@@ 5.2 Main Experiments
The performance gap between Scoring and PairS. As shown in Table 1, PairS, as a zero-shot general evaluator, achieves near state-of-the-art (SOTA) performance on all datasets, even when compared with the optimized baselines UniEval and BARTScore, which are trained on dataset-specific data. In general PairS outperforms direct scoring and G-Eval baselines in most aspects, with some exceptions, such as the consistency aspect of SummEval. We attribute this to the concentrated human scores distribution, with $86.7\%$ summaries labelled with a score of 5. These characteristics make it difficult for pairwise comparisons to yield meaningful comparisons.
@@
The Impact on different families of LLMs. PairS particularly benefits the performance of smaller LLMs. For Mistral 7B and Llama-2 7B, we observe that pairwise comparison methods generally outperform direct scoring on News Room and SummEval. Notably, Mistral 7B achieves performance comparable to GPT-3.5 on News Room. We hypothesize that this is due to their potentially similar quality of instruction tuning or RLHF stage. Llama-2 7B struggles to effectively rate summaries with scores but demonstrates better performance when comparing summaries. For HANNA, both small models do not benefit much from pairwise comparison, possibly due to the difficulty of understanding the long-context prompts or the presence of length bias. Conversely, direct scoring performs relatively well compared to summarization tasks. We believe this is because story generation evaluation does not depend on input, making the task more straightforward. For the two GPT models, PairS generally outperforms direct scoring and G-Eval baselines on all three datasets, indicating that PairS can more consistently achieve better performance for more capable models.
@@
The Benefits of Uncertain-Guided Search. PairS-beam generally exhibits higher correlations and lower Standard Error (SE), as shown in Table 3, than PairS-greedy on all datasets, indicating its ability to more robustly produce better rankings. Interestingly, the performance difference between PairS-greedy and PairS-beam is minimal for GPT-4-turbo. We attribute this to GPT-4-turbo’s better transitivity, allowing the greedy search to robustly find near-optimal rankings. There are also a few cases where PairS-beam cannot outperform PairS-greedy. This could occur when the LLM’s preference standards, the likelihood term, do not perfectly match with human annotations.
5.3 Ablation Study
@@
Transitivity. The transitivity of LLMs plays a crucial role in their effectiveness as evaluators, influencing the speed and consistency with which they achieve optimal rankings. Hence, it is a good measurement of the capability of LLMs as evaluators. We quantify transitivity through the analysis of correlation variations, specifically the standard error. We conducted 10 tuns of both PairS-greedy and PairS-beam with different random seeds. For the HANNA, we sample and shuffle the same subset of 200 stories. Table 3 shows that PairS-beam generally demonstrates better correlations and lower standard errors than PairS-greedy, which means it is more robust. We notice that GPT-3.5 consistently exhibits fewer errors across all datasets, suggesting that it can produce more consistent evaluations with itself, thus possessing better transitivity. Furthermore, different datasets show varying levels of errors, which we believe indicates the transitivity difficulty of each dataset. For example, compared to other datasets, HANNA has relatively higher errors, which means that, for the same base models, its optimal rankings are more difficult to find. This observation is expected as HANNA, being a creative story-writing dataset, has naturally more subjective human annotations, causing challenges for LLMs to align.
@@
Calibration. Unlike direct scoring, pairwise comparison has a known preference prior, which is uniform. We perform batch calibration (Zhou et al., 2023a) using this uniform human prior. Table 3 shows that calibration consistently improves performance for Mistral 7B and Llama-2 7B. However, calibration on GPT-3.5 shows relatively small improvements, likely due to its heavily skewed output probabilities (also shown in the example in Appendix 7), which makes calibration difficult and sensitive.
@@
G-3.5 54.2 35.8 37.6
- calibrated 54.7 36.5 37.8
Table 3: Calibrating LLMs. Mean of Spearman correlations (CH aspect) over 10 runs of PairS-greedy, w. and w/o calibration.
@@ 6 Conclusion
We first conduct a systematic analysis that reveals a misalignment between LLM and human evaluators, which cannot be effectively mitigated by applying calibration techniques. Inspired by RLHF, we propose that pairwise comparisons are inherently more aligned to human judgements. We formulate the evaluation as a ranking problem from the perspective of transitivity and propose PairS, an efficient, scalable, and uncertainty-guided ranking method that navigates the pairwise comparison space and estimates the MLE rankings. Extensive experimental results demonstrate that PairS is more aligned with human annotations and achieves SOTA performance compared to direct scoring and domain-optimized baselines. Furthermore, We show that PairS can be used as a measurement of LLM’s transitivity.
@@ Appendix A Additional Related Work
Calibrating LLM Evaluators. Though previous studies have shed light on using LLMs as evaluators, LLMs have been exposed to their sensitivity to prompt designs (Zhao et al., 2021; Zhou et al., 2023b), and even biased from various factors, including but not limited to positional bias (Wang et al., 2023b; Pezeshkpour & Hruschka, 2023), verbosity bias (Saito et al., 2023), and contextual bias (Zhou et al., 2023a). Calibration methods are then proposed to mitigate these biases by criteria optimization (Liu et al., 2023b), context decomposition (Li et al., 2023b), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a) and batching (Zhou et al., 2023a). However, we show that calibrations are yet insufficient in aligning LLM evaluators with humans.
Appendix B Derivations
B.1 Rank Aggregation Likelihood Derivation
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)*\prod_{i,j:\pi(i)>\pi(j)}(1-P(i\succ j)).$
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)*\prod_{i,j:\pi(i)>\pi(j)}P(j\succ i).$
@@
B.2 Transitive Ranking Likelihood Derivation
In this section, we will show how we derive the transitive ranking likelihood, Equation 4, from the non-transitive ranking likelihood, Equation 3, given the compositional stochastic transitivity assumption.
@@ B.2 Transitive Ranking Likelihood Derivation
B.3 Ranking to Scores Conversion via Quantile Matching
@@ Appendix C Datasets
Appendix D Implementation Details
D.1 Evaluation Strategies
@@
Appendix E Additional Experiments
E.1 Uncertainty threshold for different LLMs
Different base models have different logit distributions. Mistral 7B and Llama2 7B tend to have symmetric and bell-shape distributions, which means a large portion of the probabilities lie within the range of $[0.3,0.7]$, as shown in Figure 4. However, as shown in Figure 7, the probability distribution is different for GPT-3.5-turbo. Therefore, their optimal uncertainty thresholds would be different.
Figure 6: Illustration of PairS with a beam size of 2. Heatmap values show $P(A_{i}\succ B_{j})$ by GPT-3.5-turbo. The rest of settings is the same as Figure 4
@@ E.1 Uncertainty threshold for different LLMs
E.2 Theoretical and Empirical Evidence for Choice of Anchor Set Size
@@
If we take the desired confidence level as $0.8$, expected proportion of the most common category $p$ as $0.4$, error margin $e$ as $7\%$ and $deff$ as 1.5, then the sample size required for estimating five categories distribution is around 120.

### v3

@@ 5.1 Experimental Setup
Datasets. We evaluate PairS using three commonly used meta-evaluation benchmarks for two types of NLG tasks: abstractive summarization and creative story generation. Following previous research, we adopt two evaluation strategies, namely sample-level and dataset-level Spearman correlations, as detailed in Appendix  §E.1. For summarization task, we select News Room (Grusky et al., 2018) and SummEval (Fabbri et al., 2021), consistent with prior studies (Zhong et al., 2022; Fu et al., 2023). For the story generation task, we include HANNA (Chhun et al., 2022). [^3] We make the first application of pairwise preference to evaluate over thousands samples. We use an anchor size of 100 on PairS-beam-scaled in HANNA. We refer to Appendix §D for more details.
@@
Baselines. We consider two types of reference-free baselines: optimized and zero-shot evaluators. Optimized evaluators, UniEval (Zhong et al., 2022) and BARTscore (Yuan et al., 2021), mitigate the evaluation standard gaps using annotated data, making them specialized for a specific dataset or domain but not transferable. We treat them as strong but unfair metrics and use their reported scores as a reference. For zero-shot evaluators, we re-implement the following baselines: BERTScore (Zhang et al., 2020), G-Eval (Liu et al., 2023a), and GPTScore (Fu et al., 2023). We report direct scoring and G-Eval as the main baseline for fair comparisons.
@@
Table 1: LLM evaluation results. We report Spearman correlations on all datasets in terms of the following evaluation aspects: coherence (CH), fluency (FLU), relevancy (RE), informativeness (IN), consistency (CON), surprise (SU), and complexity (CX). For PairS-beam, the experiments were conducted with a beam size of 1000 and an uncertainty threshold at 0.6.
@@ 5.2 Main Experiments
Performance Gap between Direct Scoring and PairS. As shown in Table 1, PairS, as a zero-shot general evaluator, achieves state-of-the-art (SOTA) or near-SOTA performance on all datasets, even when compared with the optimized metrics UniEval and BARTScore, which are trained on dataset-specific data. In general, PairS outperforms direct scoring and G-Eval baselines in most aspects, with some exceptions, such as the consistency (CON) aspect of SummEval. We attribute this to the concentrated human scores distribution, with $86.7\%$ summaries labelled with a score of 5. These characteristics make it difficult for pairwise comparisons to yield meaningful comparisons and consequent rankings.
@@
Impact of PairS on Different (Families of) LLMs. PairS particularly benefits the performance of smaller LLMs. For Mistral 7B and Llama-2 7B, we observe that pairwise comparison methods generally outperform direct scoring on News Room and SummEval. Notably, Mistral 7B achieves performance comparable to GPT-3.5-turbo on News Room. We hypothesize that this is due to their potentially similar quality of instruction tuning. Llama-2 7B struggles to effectively rate summaries with scores but demonstrates much better performance when comparing summaries. For HANNA, both small models do not benefit much from pairwise comparison, possibly due to the difficulty of understanding the long-context prompts or the presence of length bias. Conversely, direct scoring performs relatively well compared to summarization tasks. We believe this is because story generation evaluation does not depend on input, making the task more straightforward. For the two closed-source GPT models, PairS generally outperforms direct scoring and G-Eval baselines on all three datasets, indicating that PairS can achieve better performance more consistently for more capable models.
@@
Benefits of Uncertainty-Guided Search. PairS-beam generally exhibits higher correlations and lower Standard Error (SE), as shown in Table 3, than PairS-greedy on all datasets, indicating its ability to more robustly produce better rankings. Interestingly, the performance difference between PairS-greedy and PairS-beam is minimal for GPT-4-turbo. We attribute this to GPT-4-turbo’s better transitivity, allowing the greedy search to robustly find near-optimal rankings. There are also a few cases where PairS-beam cannot outperform PairS-greedy. This could occur when the LLM’s preference standards, the likelihood term, do not perfectly match with human annotations.
5.3 Efficiency Comparison with Rank Aggregation Baselines
Figure 6: PairS-greedy and PairS-beam are more efficient than other pairwise rank aggregation methods. Both diagrams are evaluating on the coherence of SummEval. Each datapoint contains 16 summary candidates, with the number of pairwise comparisons ranging from 1 to 240. PairS-beam curves show results for beam sizes 1, 2, 5, 10, 20, 50, and 100, without uncertainty pruning. When beam size is 1, PairS-beam and PairS-greedy are equivalent.
PairS is an efficient pairwise rank aggregation method. We evaluate our method against two baseline rank aggregation techniques: 1) Win-loss rate and 2) ELO rating, which can both aggregate rankings from any number of pairwise comparisons. To assess their efficiency as aggregation algorithms, we calculate the correlations between human annotations and rankings predicted by PairS and the baselines using varying numbers of comparisons. The comparisons were random sampled without replacement. We also include the calibrated PairS, which requires double the computation by considering both permutations for each comparison pair.
Figure 6 demonstrates that PairS is more efficient than the baselines. PairS-greedy typically requires only about $30\%$ of the comparisons to achieve performance similar to ELO rating. The calibrated PairS generally outperforms the upper bound of the baselines, where all $N(N-1)$ pairwise comparions are utilized. We attribute the strong performance of PairS to its leveraging of the transitivity assumption of LLMs to identify the most informative comparison pairs, which provides a natural efficiency advantage. While win-rate and ELO aggregation methods are designed to handle systems or players with variable performance, making them more effective at dealing with changes, our task involves ranking constant response candidates on a specific aspect. In this context, sorting-based algorithms that exploit transitivity have inherent advantages.
5.4 Ablation Studies
@@
Transitivity. The transitivity assumption plays a crucial role in the effectiveness of LLMs as evaluators, impacting the speed and consistency with which they achieve rankings. Hence, it is a good measurement of the capability of LLMs as evaluators. We quantify transitivity through the analysis of correlation variations, specifically the Standard Error. We conducted 10 runs of both PairS-greedy and PairS-beam with different random seeds. For the HANNA dataset, we sample and shuffle the same subset of 200 stories. Table 3 shows that PairS-beam generally demonstrates better correlations and lower SE than PairS-greedy, which indicates that it is more robust. We notice that GPT-3.5-turbo consistently exhibits fewer errors across all datasets, suggesting that it can produce more consistent evaluations with itself, thus possessing better transitivity. Furthermore, different datasets show varying levels of errors, which we believe indicates the transitivity difficulty of each dataset. For example, compared to other datasets, HANNA has relatively higher errors, which means that, for the same base models, its optimal rankings are more difficult to find. This observation is expected as HANNA, being a creative story-writing dataset, has naturally more subjective human annotations, causing challenges for LLM alignment.
@@
Calibration. Unlike direct scoring, pairwise comparison has a known preference prior, which is uniform. We perform the same calibration method as in the previous work by Zheng et al. (2023), which averages the probabilities of both permutations of the candidates. Table 3 demonstrates that calibration consistently improves the performance of Mistral 7B and Llama-2 7B. However, calibration on GPT-3.5-turbo yields relatively small improvements, likely due to its heavily skewed output probabilities (also shown in the example in Figure 8, Appendix §F.2). The over-confident predictions make the language model difficult to calibrate.
@@
G-3.5 54.2 36.8 37.6
- calibrated 56.7 38.5 37.8
Table 3: Calibrating PairS. Mean of Spearman correlations (CH aspect) over 10 runs of PairS-greedy, w. and w/o calibration.
@@ 6 Conclusion
In this work, we have first conducted a systematic analysis of LLMs used as direct scorers for evaluation in generative tasks. The analysis has revealed a misalignment between LLM and human evaluators, which cannot be effectively mitigated by applying calibration techniques. Inspired by RLHF, we have then proposed the use of pairwise comparisons which should produce evaluation rankings that are inherently more aligned to human judgements. We have formulated the evaluation as a ranking problem from the perspective of transitivity and proposed PairS, an efficient, scalable, and uncertainty-guided ranking method that navigates the pairwise comparison space and estimates the MLE rankings. Extensive experimental results have demonstrated that PairS is more aligned with human annotations and achieves SOTA performance compared to direct scoring and domain-optimized baselines. Furthermore, we have shown that PairS can also be used to measure LLM’s transitivity.
Acknowledgements
The work is supported by the UK Research and Innovation (UKRI) Frontier Research Grant EP/Y031350/1 (the UK government’s funding guarantee for ERC Advanced Grants) awarded to Anna Korhonen at the University of Cambridge. The work has been supported in part by a Royal Society University Research Fellowship (no 221137; 2022-) awarded to Ivan Vulić, and by the UK EPSRC grant EP/T02450X/1.
@@ Appendix A Additional Related Work
Calibrating LLM Evaluators. Though previous studies have shed light on using LLMs as evaluators, LLMs have been exposed to their sensitivity to prompt designs (Zhao et al., 2021; Zhou et al., 2023), and even biased from various factors, including but not limited to positional bias (Wang et al., 2023b; Pezeshkpour & Hruschka, 2023), verbosity bias (Saito et al., 2023), and contextual bias (Zhou et al., 2024b). Calibration methods are then proposed to mitigate these biases by criteria optimization (Liu et al., 2023b), context decomposition (Li et al., 2023b), order permutation (Wang et al., 2023b; Zheng et al., 2023), ensembling (Li et al., 2023a) and batching (Zhou et al., 2024b). However, we show that calibrations are yet insufficient in aligning LLM evaluators with humans.
Appendix B Future Work
By formulating the evaluation as a ranking problem, we provide a playground for future work to develop better searching algorithms. PairS is a type of Breadth-First Search (BFS) method. However, under different transitivity assumptions, there may be a better trade-off between performance and efficiency. We encourage future work to explore and develop novel ranking algorithms that can optimize this trade-off.
Furthermore, future work can investigate alternative scaling approaches. For instance, by examining the performance of LLMs on selecting the best candidate from a set of K options, instead of performing pairwise ranking, the search efficiency could potentially be improved from $O(log_{2}N)$ to $O(log_{K}N)$. The improvement in efficiency could significantly impact the scalability and practicality of the search-based method, especially when dealing with large-scale datasets or real-world applications.
Appendix C Derivations
C.1 Rank Aggregation Likelihood Derivation
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)\times\prod_{i,j:\pi(i)>\pi(j)}(1-P(i\succ j)).$
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)\times\prod_{i,j:\pi(i)>\pi(j)}P(j\succ i).$
@@
C.2 Transitive Ranking Likelihood Derivation
In this section, we will show how we derive the transitive ranking likelihood, Equation (4), from the non-transitive ranking likelihood, Equation (3), given the compositional stochastic transitivity assumption.
@@ C.2 Transitive Ranking Likelihood Derivation
C.3 Ranking to Scores Conversion via Quantile Matching
@@ Appendix D Datasets
Appendix E Implementation Details
E.1 Evaluation Strategies
@@
Appendix F Additional Experiments
F.1 Scaling Variant Performance Loss
     Methods Mistral-7B Llama2-7B
$\rho$ Q $\rho$ Q
greedy 30.4 9,762 30.4 9,065
beam 32.0 24,733 31.1 24,006
beam-scaled 29.3 7,571 29.5 7,438
Table 4: Performance trade-off of the scaling variant. We compare the Spearman correlations ($\rho$) and numbers of the model query (# Q) for PairS-greedy, PairS-beam and PairS-beam-scaled on the Coherence aspect of the HANNA dataset. The beam size is set to 100, the anchor size is set to 100, and the uncertainty threshold is set to 0.67.
Because of the binary search in the second stage of the scaling variant, we expect there will be performance loss compared to PairS-greedy and PairS-beam. To investigate the actual trade-off, we conduct an ablation study using Mistral-7B and Llama2-7B. The results presented in Table 4 illustrate the trade-off between performance and computational cost. PairS-beam-scaled achieves comparable performance to PairS-greedy, while requiring fewer model queries. We note that the computational complexity of PairS-greedy grows at a rate of $O(N\log N)$, whereas PairS-beam-scaled will grow at a rate of $O(\log 100)$, making it more computationally efficient as the problem size increases.
F.2 Uncertainty Threshold for Different LLMs
Different base models have different logit distributions. Mistral 7B and Llama2 7B tend to have symmetric and bell-shaped distributions, which means a large portion of the probabilities lie within the range of $[0.3,0.7]$, as shown in Figure 4. However, as shown in Figure 8, the probability distribution is different for GPT-3.5-turbo. Therefore, their optimal uncertainty thresholds would be different.
Figure 7: Illustration of PairS with a beam size of 2. Heatmap values show $P(A_{i}\succ B_{j})$ by GPT-3.5-turbo. The rest of the settings are the same as Figure 4
@@ F.2 Uncertainty Threshold for Different LLMs
F.3 Theoretical and Empirical Evidence for Choice of Anchor Set Size
@@
If we take the desired confidence level as $0.8$, the expected proportion of the most common category $p$ as $0.4$, error margin $e$ as $7\%$ and $deff$ as 1.5, then the sample size required for estimating five categories distribution is around 120.
@@
Appendix G Prompt Templates
G.1 Prompts
Prompt Template Example for Pairwise Comparison
Evaluate and compare the coherence of the two following summaries. Source text: [source_text_input] Summary A: [summary_1] Summary B: [summary_2] Question: Which summary is more coherent? If the summary A is more coherent, please return A’. If the summary B is more coherent, please return B’. You must only return the choice. Answer: [output]
Table 5: A prompt template example for pairwise comparison of the coherence aspect in the summarization task.
Prompt Template Example for Direct Scoring
Evaluate the coherence of the following summary. Source text: [source_text_input] Summary: [summary] Please rate on a scale from 1 to 5, where 1 represents very low coherence and 5 indicates excellent coherence. You must only return an int score. Score: [output]
Table 6: A prompt template example for direct scoring of the coherence aspect in the summarization task.